Please show step by step. Talk to me like I am 10.

Corporation is interested in measuring the cost of each specific Hasorial ADSHEET EXERCISE as well as the weigheed average cost of capital (WACC), Weight Source of capital Long-derm debe Common stock equity5 The tax rate of the firm is currently 21%. The needed financial informa are as follows Det Nova can raise debt by selling $1,000-par-value, 6.5% cou rate, 10-year honds on which annual interest payments will be made. Toa issue, an average discount of $20 per bond needs to iven. There is n l ated flotation cost of 2% of par value. lowing terms: The Preferred stock Preferred stock can be sold under the following terms %of thepar 6% of the rity has a par value of SI00 per share, the annual dividend rate is stock is value, and the flotation cost is expected to he $4 per share. The prefcrre expected to sell for $102 before cost considerations. Common stock The current price of Novas common stock is $35 per st cash dividend is expected to be $3.25 per share next year. The firms di have grown at an annual rate of 5%, and it is expected that the dividend wil continue at this rate for the foreseeable future. The flotation costs are expected ta be approximately $2 per share. Nova can sell new common stock under th termis. Retained earnings The firm expects to have available $100,000 of retained ear ings in the coming year. Once these retained earnings are exhausted, the firm wil use new common stock as the form of common stock equity financing, (Note When measuring this cost, the firm does not concern itself with the tax beacke or hrokerage fees of owners.) ese TO DO Create a spreadsheet to answer the following questions: a. Calculate the after-tax cost of debt. b. Cakculate the cost of preferred stock. c. Calculate the cost of retained earnings. d. Calculate the cost of new common stock. e. Cakculate the firms WACC using retained earnings and the capital structure weights shown in the table above f. Calculate the firms WACC using new common stock and the capital strucrute weights shown in the table above.
Answer #1
a) Before tax cost of debt is the bonds YTM.
Inputs for YTM are annual interest $65, FV $1000,
n = 10, Price = 1000+20-20 = $960.
Using an online YTM calculator, YTM = 7.07%
After tax cost of debt = YTM*(1-t) = 7.07%*(1-21%) = 5.59%
b) Cost of preferred stock = 6/(102-4) = 6.12%
c) Cost of retained earnings using contant dividend
growth model = D1/P0+g, where D1 = next expected
dividend, P0 = Price and g = growth rate in dividends.
Substituting values,
Cost of retained earnings = 3.25/35+0.05 = 14.29%
d) Cost of new common stock = 3.25/(35-2)+0.05 = 14.85%
e) WACC = 5.59%*35%+6.12%*12%+14.29%*53% = 10.26%
f) WACC = 5.59%*35%+6.12%*12%+14.85%*53% = 10.56%